1. Bonds with longer maturities will always have lower prices than bonds pf shorter maturities with the same coupon due to time value of money. True or False
Answer #1

TRUE

bond having FV 100 aand coupon rate of 8%, Re=10%.

Bond A maturies in 5 Years

Price = coupon*PVaf(10%,5) + FV*PVif(10%,5)

Price = 8 * 3.7908 + 100 * 0.6209

Price = 92.42

Bond B matures in 1 years.

Price = 108 * (1/1.1^1) i.e. 108*0.9091 = 98.18

As in the above example Both bond have everything same except the maturity and we have seen bond A longer maturity has lower price as compared to Bond B. It is because that every time we dicount the cash flow with the multiple of 10%, decrease the cash flow and so is the lower prices.